Moses invests R3 745,35 at the end of each month at an interest rate of 14,5% per year, compounded monthly. How long will it take him to have R1 000 000,00?
1
Expert's answer
2021-07-29T12:06:34-0400

Calculating the number of months required (n):

"FV\space of\space annuity=P\times[\frac{(1+r)^n-1}{r}]"


"1,000,000=3,745.35\times[(\frac{1+\frac{0.145}{12})^n-1}{\frac{0.145}{12}}]"


"\frac{1,000,000}{3,745.35}\times0.012083=(1.012083)^n-1"


"3.226223+1==(1.012083)^n" ( now flip the equation and take log both sides)


"n\times log(1.012083)=log (4.226223)"


"n=\frac{log(4.226223)}{log(1.012083)}"


"n=119.99"


Where:

the future value of annuity = 1,000,000

the monthly payment (P) = 3,745.35

the monthly inteest rate "=\frac{0.145}{12}"


Thus, the given annuity will take 119.99 months, i.e., 120 months rounded off, or 10 years to achieve 1,000,000.00.Β